If 4 million people are collecting unemployment insurance benefits, 12 million people are officially unemployed, 113 million people are employed, and there are 2 million discouraged workers, (a) How many people are in the labor force?

(b) What is the unemployment rate?
Answer #1

(a):- total number of people in the labour force = total unemployed population + total employed population + total discouraged workers

= 12+ 113+2 = 127 million people

(b):- Unemployment rate = (total unemployed population ÷ total labour force) × 100

= (12÷127)×100
= 9.44%